if one angular bisector of a(x+1)^2+2h(x+1)(y-3)+b(y-3)^2=0 is 2x-3y+11=0 then equation of other angular bisector is

Grade:11

3 Answers

Anshuman Singh
5 years ago
See angular bisectors are always perpendicular to each other.therefore the equation of the other line should be of the form 3x+2y+c=0 and it should be satisfied by -1,3 hence the answer for c is 3
Anshuman Singh
5 years ago
Sorry -3 is the answer for c I got mistaken as the general equation of angular bisectors is x^2-y^2/a-b=xy/h. Interestingly u can observe coeff of x^2 plus coeff of y^2 =0 hence they r perpendicular
Anshuman Singh
5 years ago
I think this is the fastest method through which one can solve the given sum plz do reply if you are satisfied